My child frank pesapane goes to wiliam davies technical high school and they are not allowed to bring there books home. Frank made high honors last marking period but now is failing math .he has to match the graph with its linear system and tell how man the sulotions have for example -2 + y =1 -4x +2y=-6. Can u please help us thank sincerely Jeannette eddinger
Answer by Alan3354(69443)

he has to match the graph with its linear system and tell how man the sulotions have for example -2 + y =1 -4x +2y=-6. Can u please help us thank sincerely Jeannette eddinger
--------------------------
1st, don't put so much personal info in public, there are bad people everywhere.
=============
-2 + y =1
-4x +2y=-6
------------
Did you mean
-2x + y =1
-4x +2y=-6 ?
===============
-4x +2y = 2
-4x +2y =-6
-------------- Subtract
0 = 8
That's not true, so it's inconsistent.
--> no solution
===================
If you graph the 2 equations, you get parallel lines.
No intersection, no solution.